A car is traveling along a straight road with a velocity of 10 m/s. It begins to accelerate uniformly at time t = 0 and covers a distance of 300 m in 5 s. What is the magnitude of the acceleration?
1
Expert's answer
2019-09-05T09:10:23-0400

The position is given by:


"X_{f}=X_{o}+V_{ox}t+\frac{1}{2}a_{x}t^{2}"


Where

  • The initial position is

"X_{i}=0m"

  • The final position is

"X_{f}=300m"

  • The initial speed is

"V_{ox}=10\frac{m}{s}"

  • Acceleration

"a_{x}"

  • Time is

"t=5.00s"


Expression for acceleration

"X_{f}=X_{o}+V_{ox}t+\frac{1}{2}a_{x}t^{2}\\
X_{f}-X_{o}-V_{ox}t=\frac{1}{2}a_{x}t^{2}\\
\frac{1}{2}a_{x}t^{2}=X_{f}-X_{o}-V_{ox}t\\
a_{x}=\frac{2(X_{f}-X_{o}-V_{ox)t}}{t^{2}}"


Numerically evaluating


"a_{x}=\frac{2(300m-0m-10\frac{m}{s}*5.00s}{(5.0s)^{2}}"


FInally


"a_{x}=20\frac{m}{s^{2}}"


The acceleration of the car is

"\boxed{a_{x}=20\frac{m}{s^{2}}}"